# Jyler Menard,a and RA Mansbacha

### v2

@@ 3.1 Property prediction leads to latent space organization, even in data-sparse conditions
Fig. 4: Average best objective scores found at iteration $100$. Error bars correspond to the standard error of $M_{\text{best}}$ over five BayesOpt runs. Black dashed line separates those runs performed in the high-dimensional latent space (above the dashed line), and those runs performed in a PCA projection of the latent space (below the dashed line). Shaded regions denote the label percentage available at training time, with darker regions corresponding to a higher label percentage. Values are sorted within groups from largest to smallest. The full BayesOpt trajectories are depicted in SI Fig. S7.
Fig. 5: Average best objective scores found at iteration $500$. Error bars correspond to the standard error of $M_{\text{best}}$ over five BayesOpt runs. Black dashed line separates those runs performed in the high-dimensional latent space (above the dashed line), and those runs performed in a PCA projection of the latent space (below the dashed line). Shaded regions denote the label percentage available at training time, with darker regions corresponding to a higher label percentage. Values are sorted within groups from largest to smallest. The full BayesOpt trajectories are depicted in SI Fig. S7.
@@ 3.4 Searching through a PCA projection is associated with more exploration
To further understand BayesOpt performance in our spaces, we examine how much exploration occurs; that is, how much of the space is probed by the search. We quantify the amount of exploration that occurs throughout a BayesOpt run through four lenses: (1) the amount of the total latent space (input) a BayesOpt run spans, which we estimate using the hypervolume of the sampled points; (2) the amount of the scores (output) a BayesOpt run spans, which we estimate using the variance in objective scores sampled and (3) the distance from the oracle SVR’s training set; and (4) the amount of the relevant latent space (relevant input) a BayesOpt run spans, which we estimate using, which we estimate the total path length of the sampled points in two-dimensional projections of the space. We first characterize the exploration and then assess its relationship with exploitation as previously quantified by achieved best scores.
We compute the hypervolume of the input points sampled in each BayesOpt run, which is essentially a measure of the extent to which we explore different objective function inputs. The hypervolume was computed using the pygmo python package 55; it computes the hypervolume of a set of points compared to a reference point; to choose our reference point, we used the corner of the search box with maximum value in each dimension ($10$ in each dimension). When looking at the distribution of resulting hypervolumes, we observe more hypervolume is explored on average when optimizing in a PCA search space compared to the high-dimensional search space; the PCA distribution (Fig. 6a orange) is shifted to the right of the high-dimensional search space distribution (Fig. 6a, blue), with $\mu_{PCA,HV}=1.1\cdot 10^{-17}>0.55\cdot 10^{-17}=\mu_{Id,HV}$. This suggests that a broader slice (about 50% more) of the underlying high-dimensional latent space was explored despite the PCA runs being confined to a subspace of it.
@@
Fig. 6: Distributions of exploration quantities for optimization done in the full latent space ("Id", blue) or the PCA projection of it (PCA, orange). (a) Distribution of each run’s computed hypervolume. (b) Distribution of each BayesOpt run’s variance in scores sampled, $\text{Var}(M)$. (c) Distribution of each run’s best sampled point’s distance from the oracle’s training set, computed using Eqn. 8. (d) Distribution of each run’s total path length, computed using Eqn. 9. For (c) and (d) the distances are computed in the two PCs most correlated with the oracle values.
@@
We further examined the BayesOpt runs in their associated latent spaces qualitatively. To do this, we plotted the peptides sampled during a BayesOpt run in a two-dimensional PCA projection of the associated latent space using the two PCs most correlated to oracle values. In Figure 7, we depict a full example trajectory showing the latent space (panel a), the structures of 17 selected peptides along the run as predicted by ESMFold (panel b), and their corresponding scores (panel c). This trajectory illustrates the best-scoring run of five that were performed in a 20-dimensional projection of the oracle-organized latent space. This figure has a clear interpretation and allows us to gain a strong physical intuition for the progress of the search. We see the BayesOpt trajectory begin at a low objective value (1) and rapidly increase by moving to the right in the latent space, towards the average direction of increased objective value (lower MIC). After attaining point 7, BayesOpt performs a large explorative jump to the left to point 8, which only marginally increases the best score. It then proceeds by generally drifting to the right and eventually jumping upwards (point 16). Concomitantly, we see that as the search proceeds, the predicted structures increase strongly in helicity. This is a reasonable proxy for antimicrobialness, particularly given the over-representation of $\alpha$-helical AMPs in most datasets; however, it does point to a limitation of the oracle model in place of true experimental data. This provides another compelling reason for the use of low label percentage highly-relevant information rather than “filling in the gaps” with less relevant information or predictors.
@@
In Fig. 8, we visualize the best-scoring run of five for all oracle-organized latent spaces at 100% and 2% label percentage. In the Supplementary Information, we also visualize the best run in each physicochemical property–organized space and label percentage, coloured by oracle values (SI Figs S11, S12, S13, S14). As previously quantified, we visually observe that BayesOpt runs tend to cover more of the 2D visualizations when their associated search space was a PCA projection (e.g., compare Fig. 8g with Fig. 8h-l, the rows of SI Fig. S12, or the rows of SI Fig. S15). Additionally, there is substantially greater spatial exploration when the percentage of labels is low, although as we have seen, BayesOpt is still able to converge to satisfactory optima in this case. In particular, when comparing the low-label regime (2%) with the full label regime (100%), we observe that the BayesOpt runs explore substantially more of the visualization; this is especially the case when BayesOpt runs search through the top two, five, or ten PCA dimensions.
@@
Fig. 7: Best peptides identified during a BayesOpt run. (a) PCA plot of the associated oracle-organized latent space (at 2% of labels available), with barplot depicting an average decrease in predicted-$\log_{10}(\text{MIC})$ from left-to-right across the latent space. The latent space locations of the best peptides are depicted with large circle markers; white integers are the sequence in which these peptides were identified during a BayesOpt run, but are not the iteration at which they were identified. (b) ESMFold-predicted structures 56 of the best peptides found during the BayesOpt run, visualized using ChimeraX 57. In this case we note a strong increase in helicity with oracle value, a reasonable proxy that the model might learn for antimicrobialness, although it suggests a limitation of the oracle for actual production, which would be ameliorated by the use of the underlying information in its stead. (c) The corresponding best score $M_{\text{best}}$ found throughout the BayesOpt run.
@@ 4 Discussion and Conclusions
In this article, we have trained a series of Transformer-based VAE models for peptide sequence generation, which we have used as continuous search spaces for latent Bayesian optimization. We presented a new version of latent Bayesian optimization and compared it with traditional latent BayesOpt. Inspired by a desire to better align visualization/analysis techniques with algorithmic properties, we took PCA projections of the latent spaces–previously used solely for visualization–and performed BayesOpt directly in them. We assessed the properties of the latent spaces themselves, compared our low-dimensional BayesOpt search with direct Bayesian Optimization in the full 64-dimensional latent space of the trained generative model, and connected latent space properties with resultant properties of the BayesOpt search trajectories.
In line with previous work 22, 21 we found that jointly-training a property-predictor and a VAE can lead to an organized latent space. Expanding on that work, we investigated whether such organization persists even in a semi-supervised scenario, finding evidence that just $2\%$ of property labels suffices to induce organization along that property; additionally we showed that jointly-training with multiple properties can lead to organization along each property.
@@
Despite the fact that the performance of projected latent BayesOpt can be more variable and more sensitive to latent space organization than latent BayesOpt, projected latent BayesOpt has a significant advantage in terms of interpretability. We can more easily and accurately visualize and understand the trajectories of projected latent BayesOpt. In addition, we showed that projected latent BayesOpt explores more of the search space than latent BayesOpt, in terms of both full hypervolume and exploration along relevant dimensions. Furthermore, the variability in performance may also be framed as heightened exploration of output scores sampled, which we showed to be correlated with better performance. Although projected latent BayesOpt has a weakness, we suggest that it can be ameliorated by running multiple short trajectories in spaces of different dimensionalities, and it comes with the advantage of better understanding of the search procedure. We also note that this work further underscores an important point we have previously raised 21, which is that naive use of low-dimensional projection algorithms to visualize latent spaces of generative models can be misleading. The community should not over-rely on the assumption that a two-dimensional PCA or t-SNE projection is sufficient to reveal qualities of the latent space, at least not without further analysis.
In sum, latent Bayesian optimization can provide an approach to identifying highly-potent antimicrobial peptides leveraging the continuous-valued latent spaces of generative models. We found that VAEs based on transformers can have their latent spaces organized by multiple different physicochemical properties simultaneously through the joint-training of multiple property predictors with the VAE. We further compared whether using physicochemical properties to organize the latent space is better than using few oracle values, finding evidence that using few oracle evaluations can be better when restricting the search space to linear projection of the latent space; i.e. having a more relevant organizing property with few labels may be a good use case for searching through PCA projection of the latent space. Using the linear projection comes with the additional benefit of to being more interpretable, and easier to visualize.
@@ Acknowledgements
  • Arnold et al. 2025 A. Arnold, S. McLellan and J. M. Stokes, npj Antimicrobials and Resistance, 2025, 3, 18.
@@
  • 35 M. Thoma and J. Hahnfeld, ProPy3: Python3 package for computing protein descriptors., https://pypi.org/project/propy3/.
@@
  • 58 https://www.comet.com/site/.
  • 59 https://www.hyundai.com/worldwide/en/eco/ioniq6/technology.
Appendix S1 Dataset
To train our SVR model, we used the Giant Repository of AMP Activities (GRAMPA) dataset published in Witten and Witten 32. This dataset contains peptide sequences and their associated Minimum Inhibitory Concentration (MIC).
Fig. S1: Distribution of peptide sequence lengths in our dataset.
Appendix S2 SVR Oracle
As our oracle of antimicrobial activity, we developed a Support Vector Regression model predicting the base-10 logarithm of MIC, $\log_{10}(\text{MIC})$. In this section, we provide additional details about the dataset, feature selection method, and performance of the SVR. We opted not to include this information in the main text as it was not immediately relevant to the Bayesian Optimization questions we were asking.
@@
Table S1: Number of features at each stage during building of the SVR oracle.
Appendix S3 Energy consumption
We estimate the total energy consumption of training the deep learning models to 100 epochs. The system we performed training on was a NVIDIA V100 Volta (32G HBM2 memory) through the Digital Research Alliance of Canada’s Cedar cluster. From comet-ml, we observed an average power usage of $\approx 275W$ throughout a given training job 58. To train a given model to 100 epochs, approximately 2424hrs were required, yielding $275\cdot 24/1000=6.6$ kWh per 100 epochs. For the main text, we trained 27 models to 100 epochs, giving an approximate total energy usage of 178.2178.2kWh.
The Hyundai Ioniq 6 is a 2022 battery electric sedan. Its long-range battery capacity is 77.477.4kWh, corresponding to an estimated range of 614614km 59, giving $\approx 0.126$kWh/km. Comparing to our total energy usage, we obtain an approximate car range using energy from 100 epochs as $\approx 52.36$km. Therefore, the total approximate range of a Hyundai Ioniq 6 using the energy from training our main text models is 1413.631413.63km, or we could fully recharge $2.3$ Hyundai Ioniq 6 sedans.
System NVIDIA V100 Volta
Average GPU energy utilization (Watts) 275
Approx. hours for 100 epochs 24
kWh per 100 epochs 6.6
Hyundai Ioniq 6 battery capacity 77.4 kWh
Hyundai Ioniq 6 estimated range 614 km
Energy per km 0.1261 kWh/km
Approx. car range from 100 epochs 52.36 km
Number of 100 epoch runs 27
Approximate total energy used 178.2 kWh
Approximate equivalent car range 1413.63 km
Appendix S4 Additional Model Analysis
@@
Table S3: Principal component pairs used for visualization in Fig. 1. The location of each pair corresponds to their subpanel.
@@
S4.1 Summary of models
S4.2 Relation between organizing properties and predictive oracle
@@
Boman Charge Hydrophobicity

### v3

@@ 3.1 Property prediction leads to latent space organization, even in data-sparse conditions
Fig. 5: Average best objective scores found after (a) $100$ iterations and after (b) $500$ iterations. Error bars correspond to the standard error of $M_{\text{best}}$ over five BayesOpt runs. Black dashed line separates those runs performed in the high-dimensional latent space (above the dashed line), and those runs performed in a PCA projection of the latent space (below the dashed line). Shaded regions denote the label percentage available at VAE training time, with darker regions corresponding to a higher label percentage. In (a) values are sorted within groups from largest to smallest, while (b) uses the same ordering as (a). Each class refers to the properties used for organizing the latent space, with "VAE" referring to an unorganized latent space. The full BayesOpt trajectories are depicted in SI Fig. S7.
3.4 Bayesian optimization in a linearly-projected space may outperform that in a nonlinearly projected space
We additionally compared performing BayesOpt in a linearly-projected space to performing BayesOpt in a nonlinearly-project lower dimensional space on the fly by employing Gaussian Process Deep Kernel Learning (GP-DKL) 55. GP-DKL is a variant of a Gaussian Process that uses a neural network to augment the kernel function. By learning a projection $g_{\theta}(\vec{x})$ mapping design points into a feature space that the kernel function operates on, the conventional BayesOpt kernel, $k(\vec{x}_{1},\vec{x}_{2})$, is replaced by $k(g_{\theta}(\vec{x}_{1}),\;g_{\theta}(\vec{x}_{2}))$, where $g_{\theta}$ is a neural network. The weights $\theta$ of $g_{\theta}$ are fit simultaneously with the parameters of the kernel $k(,)$ at each BayesOpt iteration. GP-DKL is an a promising alternative approach to deriving a low-dimensional space for BayesOpt to traverse, as it operates directly on the high-dimensional space and leverages a learnable non-linear projection (as opposed to the static, linear one we have used) enabling the GP to determine the projection necessary to find an optimal point.
Here we implemented GP-DKL, and performed BayesOpt with five-dimensional GP-DKL in five of our TransVAE spaces: unorganized, BCH-organized with either 100% or 2% of the labels, and oracle-organized with either 100% or 2% of the labels. Across three different architectures for the GP-DKL neural network (dkl-v1, dkl-v2, dkl-v3; see SI Sect. S4.6 for more details) we find that, in all but one case, BayesOpt in a five-dimensional PCA-projected space is able to consistently perform as well as, or better than, GP-DKL (Figs 6 and S13 ).
In an unorganized TransVAE model, two architectures from GP-DKL (dkl-v2, orange; dkl-v3, green) perform as well BayesOpt in a PCA-projection (red), but one GP-DKL variant (dkl-v1, blue) does slightly better than searching in the PCA-projection. However, we observe that searching directly in the high-dimensional space (purple) results in substantially higher values attained than either GP-DKL or BayesOpt in the linear-projection, thus pointing to fundamental limitations of searching in a lower-dimensional space without additional information to permit an informed choice of directions.
Once we move to organized spaces, there is a clearer difference. In the BCH-organized spaces, all three GP-DKL variants have average best scores below zero while the BayesOpt searching through the high-dimensional space and searching through the PCA-projection are able to by at least the 100th iteration. This trend also holds when we look at the BayesOpt runs done in the oracle-organized spaces. This may be a result of differences in data availability to the different algorithms. The PCA projector is fit to the (embedded) training set of the VAE, which is O(100k) points. This projector is then not updated throughout the BayesOpt loops. In DKL, the neural network projector is fit during the BayesOpt iterations on the fly, meaning it can search in different directions while severely restricting the amount of data available ( O(100) points in our scenario ) compared to the PCA projection. Further work could be done to optimize a nonlinear projection scenario.
Fig. 6: Average best objective scores found after $500$ iterations. Error bars correspond to the standard error of $M_{\text{best}}$ over five BayesOpt runs. Abbreviations correspond to neural network architectures described in SI Sect. S4.6. The full BayesOpt trajectories are depicted in SI Fig. S7.
@@ 3.5 Searching through a PCA projection is associated with more exploration
To further understand BayesOpt performance in our spaces, we examine how much exploration occurs; that is, how much of the space is probed by the search. We quantify the amount of exploration that occurs throughout a BayesOpt run through four lenses: (1) the amount of the total latent space (input) a BayesOpt run spans, which we estimate using the hypervolume of the sampled points; (2) the amount of the scores (output) a BayesOpt run spans, which we estimate using the variance in objective scores sampled and (3) the distance from the oracle SVR’s training set; and (4) the amount of the relevant latent space (relevant input) a BayesOpt run spans, which we estimate using, which we estimate the total path length of the sampled points. We first characterize the exploration and then assess its relationship with exploitation as previously quantified by achieved best scores.
We compute the hypervolume of the input points sampled in each BayesOpt run, which is essentially a measure of the extent to which we explore different objective function inputs. The hypervolume was computed using the pygmo python package 56; it computes the hypervolume of a set of points compared to a reference point; to choose our reference point, we used the corner of the search box with maximum value in each dimension ($10$ in each dimension). When looking at the distribution of resulting hypervolumes, we observe more hypervolume is explored on average when optimizing in a PCA search space compared to the high-dimensional search space; the PCA distribution (Fig. 7a orange) is shifted to the right of the high-dimensional search space distribution (Fig. 7a, blue), with $\mu_{PCA,HV}=1.1\cdot 10^{-17}>0.55\cdot 10^{-17}=\mu_{Id,HV}$. This suggests that a broader slice (about 50% more) of the underlying high-dimensional latent space was explored despite the PCA runs being confined to a subspace of it.
@@
Fig. 7: Distributions of exploration quantities for optimization done in the full latent space (“Id”, blue) or the PCA projection of it (PCA, orange). Distribution of each run’s: (a) computed hypervolume; (b) variance in scores sampled, $\text{Var}(M)$; (c) best sampled point’s distance from the oracle’s training set (Eqn. 8); (d) total path length (Eqn. 9). For (c) and (d) the distances are computed in the two PCs most correlated with the oracle values.
@@
We further examined the BayesOpt runs in their associated latent spaces qualitatively. To do this, we plotted the peptides sampled during a BayesOpt run in a two-dimensional PCA projection of the associated latent space using the two PCs most correlated to oracle values. In Figure 8, we depict a full example trajectory showing the latent space (panel a), the structures of 17 selected peptides along the run as predicted by ESMFold (panel b), and their corresponding scores (panel c). This trajectory illustrates the best-scoring run of five that were performed in a 20-dimensional projection of the oracle-organized latent space. This figure has a clear interpretation and allows us to gain a strong physical intuition for the progress of the search. We see the BayesOpt trajectory begin at a low objective value (1) and rapidly increase by moving to the right in the latent space, towards the average direction of increased objective value (lower MIC). After attaining point 7, BayesOpt performs a large explorative jump to the left to point 8, which only marginally increases the best score. It then proceeds by generally drifting to the right and eventually jumping upwards (point 16). Concomitantly, we see that as the search proceeds, the predicted structures increase strongly in helicity. Given the over-representation of $\alpha$-helical AMPs in most datasets, helicity is a reasonable proxy for antimicrobialness that the BayesOpt is exploiting; a straightforward example of reward hacking. This confirmation of reward hacking provides another reason for using the most relevant data possible (less hackable), even if less of it is available (because e.g. it is expensive to compute).
@@
In Fig. 9, we visualize the best-scoring run of five for all oracle-organized latent spaces at 100% and 2% label percentage. (For more information cf. also SI Figs S15-S18). As previously quantified, we visually observe that BayesOpt runs tend to cover more of the 2D visualizations when their associated search space was a PCA projection (e.g., compare Fig. 9g with Fig. 9h-l ). Additionally, there is substantially greater spatial exploration when the percentage of labels is low, although as we have seen, BayesOpt is still able to converge to satisfactory optima in this case. In particular, when comparing the low-label regime (2%) with the full label regime (100%), we observe that the BayesOpt runs explore substantially more of the visualization; this is especially the case when BayesOpt runs search through the top two, five, or ten PCA dimensions.
@@
Fig. 8: Best peptides identified during a BayesOpt run. (a) PCA plot of the associated oracle-organized latent space (at 2% of labels available), with barplot depicting an average decrease in predicted-$\log_{10}(\text{MIC})$ from left-to-right across the latent space. The latent space locations of the best peptides are depicted with large circle markers; white integers are the sequence in which these peptides were identified during a BayesOpt run, but are not the iteration at which they were identified. (b) ESMFold-predicted structures 57 of the best peptides found during the BayesOpt run, visualized using ChimeraX 58.In this case we note a strong increase in helicity with oracle value, a reasonable – but biophysically misleading – proxy that the oracle might learn for antimicrobialness, which may be ameliorated with higher-fidelity evaluations such as simulation. (c) The corresponding best score $M_{\text{best}}$ found throughout the BayesOpt run.
@@ 4 Discussion and Conclusions
In this article, we have trained a series of Transformer-based VAE models for peptide sequence generation, which we have used as continuous search spaces for latent Bayesian optimization. We presented a version of latent Bayesian optimization with additional dimensionality reduction, and compared it with traditional latent BayesOpt and BayesOpt with Deep Kernel Learning. Inspired by a desire to better align visualization/analysis techniques with algorithmic properties, we took PCA projections of the latent spaces–previously used solely for visualization–and performed BayesOpt directly in them. We assessed the properties of the latent spaces themselves, compared our low-dimensional BayesOpt search with direct Bayesian Optimization in the full 64-dimensional latent space of the trained generative model, and connected latent space properties with resultant properties of the BayesOpt search trajectories.
In line with previous work 26, 25 we found that jointly-training a property-predictor and a VAE can lead to an organized latent space. Expanding on that work, we investigated whether such organization persists even in a semi-supervised scenario, finding evidence that just $2\%$ of property labels suffices to induce organization along that property; additionally we showed that jointly-training with multiple properties can lead to organization along each property.
@@
Despite the fact that the performance of projected latent BayesOpt can be more variable and more sensitive to latent space organization than latent BayesOpt, projected latent BayesOpt has a significant advantage in terms of interpretability. We can more easily and accurately visualize and understand the trajectories of projected latent BayesOpt. In addition, we showed that projected latent BayesOpt explores more of the search space than latent BayesOpt, in terms of both full hypervolume and exploration along relevant dimensions. Furthermore, the variability in performance may also be framed as heightened exploration of output scores sampled, which we showed to be correlated with better performance. Although projected latent BayesOpt has a weakness, we suggest that it can be ameliorated by running multiple short trajectories in spaces of different dimensionalities, and it comes with the advantage of better understanding of the search procedure. We also note that this work further underscores an important point we have previously raised 25, which is that naive use of low-dimensional projection algorithms to visualize latent spaces of generative models can be misleading. The community should not over-rely on the assumption that a two-dimensional PCA or t-SNE projection is sufficient to reveal qualities of the latent space, at least not without further analysis.
In sum, latent Bayesian optimization can provide an approach to identifying highly-potent antimicrobial peptides leveraging the continuous-valued latent spaces of generative models. We found that VAEs based on transformers can have their latent spaces organized by multiple different physicochemical properties simultaneously through the joint-training of multiple property predictors with the VAE. We further compared whether using physicochemical properties to organize the latent space is better than using few oracle values, finding evidence (Sect. 3.3) that using few oracle evaluations can be better when restricting the search space to linear projection of the latent space; i.e. having a more relevant organizing property with few labels may be a good use case for searching through PCA projection of the latent space. Using the linear projection comes with the additional benefit of to being more interpretable, and easier to visualize.
Although we have primarily discussed our work in the context of AMP design, given that BayesOpt is an efficient, iterative algorithm for searching a design space to optimize a function under limited data, our results extend straightforwardly to any peptide design question. The only limitation is the definition of a relevant objective function. For example, one could optimize for cell-penetrating peptides through in silico measurements of maximum force experienced by a peptide as it is pulled through a lipid bilayer in constant-velocity steered molecular dynamics; alternatively, one could optimize for anticancer peptides through in vitro measurements of the minimum peptide concentration required to inhibit cancer cell growth, or through in silico approximations of binding against a target protein. More broadly, the application of BayesOpt to a particular peptide sequence design challenge is limited primarily by the availability of an oracle that can be regularly queried for estimates of the underlying objective.
@@
Data Availability Statement
The code for model training and analysis can be found at https://github.com/Mansbach-Lab/compare-latent-spaces-amps/tree/main. We note that for certain model training we employed the publicly available GRAMPA dataset 36. Trained model checkpoints and datasets are released through Zenodo https://zenodo.org/records/17872434 with DOI https://doi.org/10.5281/zenodo.17872449.
@@ Acknowledgements
  • Fjell et al. 2007 C. D. Fjell, R. E. Hancock and A. Cherkasov, Bioinformatics, 2007, 23, 1148–1155.
  • Fjell et al. 2009 C. D. Fjell, H. Jenssen, K. Hilpert, W. A. Cheung, N. Panté, R. E. Hancock and A. Cherkasov, Journal of medicinal chemistry, 2009, 52, 2006–2015.
  • Randou et al. 2013 E. G. Randou, D. Veltri and A. Shehu, 2013 IEEE 3rd International Conference on Computational Advances in Bio and medical Sciences (ICCABS), 2013, pp. 1–6.
  • Lee et al. 2017 E. Y. Lee, M. W. Lee, B. M. Fulan, A. L. Ferguson and G. C. Wong, Interface focus, 2017, 7, .
  • Spänig and Heider 2019 S. Spänig and D. Heider, BioData Mining, 2019, 12, 7.
@@
  • Boman 2003 H. G. Boman, Journal of internal medicine, 2003, 254, 197–215.
  • Kyte and Doolittle 1982 J. Kyte and R. F. Doolittle, Journal of molecular biology, 1982, 157, 105–132.
@@
  • Wilson et al. 2016 A. G. Wilson, Z. Hu, R. Salakhutdinov and E. P. Xing, Proceedings of the 19th International Conference on Artificial Intelligence and Statistics, 2016, pp. 370–378.
@@
  • 61 M. Thoma and J. Hahnfeld, ProPy3: Python3 package for computing protein descriptors., https://pypi.org/project/propy3/.
  • Kingma and Ba 2017 D. P. Kingma and J. Ba, Adam: A Method for Stochastic Optimization, 2017, http://arxiv.org/abs/1412.6980, arXiv:1412.6980 [cs].
  • 66 https://www.comet.com/site/.
  • 67 https://www.hyundai.com/worldwide/en/eco/ioniq6/technology.
Appendix S1 Supplementary information for Towards best practices in low-dimensional semi- supervised latent Bayesian optimization for the design of antimicrobial peptides
Appendix S2 Support Vector Regression Oracle
To serve as an oracle to rapidly evaluate our “true” objective function, we trained a Support Vector Regression (SVR) model on data compiled by Witten & Witten (2019)36 by using the scikit-learn SVR implementation 59. This dataset contains peptide sequences and their associated Minimum Inhibitory Concentration (MIC), which we split into a training (4145 sequences) and a validation (2615 sequences) set. This dataset primarily contains sequences of lengths <50<50 amino acids.
@@
For the outputs of the oracle, we employed the base-10 logarithm of the MIC values $\log_{10}(\text{MIC})$ to ensure the values are distributed more symmetrically (Fig. S1). For the inputs to the oracle, we used easily-computed physico-chemical descriptors. We used a previously-validated feature selection procedure60 to identify the most relevant subset from a larger set of possible descriptors.
Computing all descriptors available in the ProPy3 Python package 61 resulted in 1529 features per sequence (Table S1). We reduce the number of features necessary for the model via a two-step feature selection procedure. We perform an initial feature screening stage using the minimum Redundancy Maximum Relevance (mRMR) algorithm62, in which one attempts to identify those features that have the least overlap that are the most relevant to the prediction of the output of interest. We defined redundancy as the mutual information between pairs of physico-chemical features and relevancy as the mutual information between features and predicted outputs (i.e. $\log_{10}(\text{MIC})$).
Abbreviation Name Description
AAC Amino Acid Composition
The fraction of a sequence composed of
a given amino acid.
DPC Dipeptide Composition
The fraction of a sequence composed of
a given pair of amino acids.
MBauto
Normalized Moreau-Broto
Autocorrelation
Normalized autocorrelation of properties,
where property similarities are computed.
Moranauto Moran autocorrelation
Autocorrelation of properties, where deviations
from the average are computed.
Gearyauto Geary Autocorrelation
The squared-difference of a property at different
sequence locations is computed, and normalized
by the variance.
CTD Composition, Transition, Distribution
Amino acids are grouped based on 7 different
properties. Then for each property the
composition of the sequence, the number of
transition between groups, and the distribution
of groups is computed.
SOCN Sequence Order Coupling Numbers
Quantifies the extent to which amino acids
are ’coupled’ in a sequence based on amino
acid physicochemical distances.
QSO Quasi Sequence Order
Composition and amino acid correlation,
computed using amino acid
physicochemical "distance" matrices.
PAAC Pseudo Amino Acid Composition
Computes the amino acid composition of a
sequence while maintaining a notion of how
correlated amino acids’ hydrophilicity,
hydrophobicity, and mass are
throughout the sequence.
Table S1: Physicochemical descriptors computed using ProPy3 python package61.
[…]
Rank Name
LASSO
Weight
Brief Description
1 tausw2 -0.19 Quasi Sequence Order coupling number
2 PolarizabilityC3 -0.14 Polarizability Composition
3 ChargeT12 -0.13 Number of charge transition between charged and neutral AAs
4 NormalizedVDWVC2 -0.09 van der waals volume
5 PAAC20 0.08
Type I Pseudo amino acid composition descriptors;
it uses hydrophobicity, hydrophilicity, residue mass.
6 QSOSW39 -0.07 Quasi Sequence Order using Schneider-Wrede distance matrix
7 PAAC21 0.06 Type I Pseudo amino acid composition descriptors
8 MoreauBrotoAuto_Steric2 0.06 Moreau-Broto Autocorrelation
9 ChargeD2100 0.05 Charge distribution group 2 100%
10 MoranAuto_Hydrophobicity5 0.05 Moran Autocorrelation
Table S2: The top-10 features selected after mRMR and LASSO. Note that this is a subset of the 149 features we kept after our feature selection procedure.
@@
Table S3: Number of features at each stage of feature selection when building the SVR oracle.
Appendix S3 Additional TransVAE information and analysis
S3.1 Additional model details
We trained variational autoencoders (VAEs) with transformer architectures based on Renaud and A. Mansbach 25. Model architecture was chosen based on work done in Renaud and A. Mansbach 25, with specific hyper-parameter values and descriptions listed in Table S4. To optimize the weights of the neural networks, we used the Noam optimizer40, a variant of the Adam optimizer 65 with a varying learning rate. The output dimensions of the property predictor module varied over none, 1, 2, 3, depending on the number of organizing properties. All models were trained to 100 epochs, with the total training loss <0.9<0.9 and validation losses <0.95<0.95, with most models have a total validation loss <0.8<0.8 (Fig. S3).
Hyperparameter Value Description
$d_{\text{latent}}$ 64 Dimensionality of latent space
$d_{\text{model}}$ 512 Dimensionality of model
$N_{\text{block}}$ 3 Number of Encoder/Decoder blocks
$d_{\text{ff}}$ 256 Dimensionality of feedforward layers
$h$ 4 Number of attention heads per multi-head attention layer
$d_{\text{pp}}$ 64 Dimensionality of property predictor layers
$\text{depth}_{\text{pp}}$ 2 Number of property predictor layers
$d_{\text{pp,out}}$ 1,2,3 Output dimensionality of property predictor
$\beta_{\text{pp},0}$ 0 Initial value for property predictor annealing coefficient
$\beta_{\text{pp}}$ 1 Final value for property predictor annealing coefficient
$\beta_{\text{KL},0}$ 0 Initial value for KL annealing coefficient
$\beta_{\text{KL}}$ 0.05 Final value for KL annealing coefficient
$N_{\text{batch}}$ 256 Batch size during training
$\beta_{1}$ 0.9 Adam optimizer’s $\beta_{1}$
$\beta_{2}$ 0.98 Adam optimizer’s $\beta_{2}$
lr $(d_{model})^{-0.5}\cdot\min\Big(\text{step}^{-0.5},\text{step}\cdot(n_{\text{warmup}})^{-1.5}\Big)$ Learning rate set adaptively using the Noam optimizer
$n_{\text{warmup}}$ 10000 Number of warmup steps for the Noam optimizer
$p_{\text{dropout}}$ 0.1 Dropout rate
Table S4: Hyper-parameters for the model architecture and training. The TransVAE consists of an encoder and decoder module, each consisting of $N_{\text{block}}$ blocks of multi-head self-attention feeding into feedforward layers.
Fig. S2: Distribution of peptide sequence lengths in our dataset.
S3.2 Summary of models
S3.3 Additional PCA projection analysis
Principal Component Analysis (PCA) determines a set of ordered basis vectors for a set of data points, where each basis vector in the ordering captures a decreasing amount of variance in the original cloud of data points. In the main text, we frequently use only the top five principal components as a lower dimensional representation of a full latent space. Here we provide the full distribution of principal components (Fig. S4). We observe that the top five principal components capture between 17% and 22% of the variance in the VAE training set.
Along with explained variance, we calculate the correlation between each of the top five principal components, and the organizing property (or each individual property in the case of multi-property organization), allowing us to best visualize the property-induced organization of the latent space. In Fig. 2, we visualized the extent to which the latent space is organized by Boman index, charge, and hydrophobicity, when the underlying TransVAE model was being organized by all three simultaneously. Table S5 provides the PC pair used for visualization. Charge is frequently most organized along (most correlated with) PCs separate from Boman and Hydrophobicity. Boman and Hydrophobicity frequently have overlapping, but not always the same, PC pair; while Boman and hydrophobicity have similar trends, they are not quite the same, this can suggest that the 2D projection, and the linear correlation used to select PCs, is hiding differences in organization.
@@
Table S5: Principal component pairs used for visualization in Fig. 2. The location of each pair corresponds to their subpanel in Fig. 2.
S3.4 Relation between organizing properties and predictive oracle
We investigated in some detail the correlations between the principal components and the “bridge variables,” that is different physicochemical properties, including organizing variables. We observed that sequence length was most correlated with the first principal component across 21/27 of the models (SI Fig S5g), with the second or third principal component being correlated with the property used during training to organize the latent space (SI Fig. S5h for Boman index; SI Fig. S5i for charge; SI Fig. S5j for hydrophobicity; SI Fig. S5k for oracle). Although we may desire the first principal component to be most correlated with the organizing property in most cases, we note that the top five principal components have similar amounts of explained variance, and that they capture on the order of $20\%$ of the total variance (SI Fig. S4). Additionally, previous work found sequence length can be an easy-to-learn pattern25. Patterns that are straightforward to recognize may become unexpectedly strongly correlated with a PC direction.
@@
Boman Index Charge Hydrophobicity